OpenStudy (lovemygolden128):

what is the product of (3 + 7i)(1 − i), and write the answer in standard form.

OpenStudy (_citrus_):

\[10 + 4i\] That is your answer it is already in standard form :)

OpenStudy (_citrus_):

You can solve it by using the FOIL method, then combine the real and imaginary parts of the expression

OpenStudy (unklerhaukus):

(3 + 7i)(1 - i) = 3 - 3i +7i -7i^2 = 3 + (-3+7)i + 7